Find the maximum speed vmax of the photoelectrons leaving the surface of silver

under the influence of the ultraviolet radiation of the wave length λ = 0.155μm. The work

function of silver is A = 4.7 eV.


1
Expert's answer
2022-01-10T13:59:05-0500

"\lambda=0.155\mu m"

A=4.7eV

We know that

"E=\frac{hc}{\lambda}"

"h=6.625\times10^{-34}Js"

"E=\frac{6.625\times10^{-34}\times3\times10^8}{1550\times1.6\times10^{-19}}"

"E=\frac{12420}{1550}eV"

"E=8eV"

We know that

Photo electric effect

"E-A=eV_s"

"(8-4.7)eV=eV_s"

"eV_s=3.3eV"

We know that

"\frac{1}{2}mv^2=eV_s"

"v=\sqrt\frac{2eV_s}{m}"


"v=\sqrt{\frac{2\times3.3\times1.6 \times10^{-19}}{13\times9.1\times10^{-31}}}=2.98\times10^{5}m/sec"